# Analysis of Internet of Things implementation barriers in the cold supply chain: an integrated ISM-MICMAC and DEMATEL approach

### v2

@@ 2.3 IoT implementation barriers
IoT deployment is challenging due to device and data security, interoperability issues, scalability, trust, privacy concerns, financial issues, a lack of qualified personnel, and compliance with data privacy laws. First, it is necessary to identify the barriers that hinder IoT implementation in the CSC to overcome these challenges. Table 1 lists the barriers to implementing IoT in CSC.
@@
The CSC has strict temperature and data management guidelines. It may be challenging to implement IoT technology and meet these requirements. In the pharmaceutical and food industries, strict regulations are tough to follow. Failing to satisfy these regulations may result in product recalls or legal concerns.
[39] [40] [41] [42]
02
High implementation cost
@@
In order to adopt the IoT, it may be necessary to make a sizable initial investment in things like sensors, devices, and infrastructure. Transporting and storing cold requires plenty of energy. Costs may rise with power prices.
[43] [28] [29] [38]
03
Lack of skilled workforce
@@
Multiple sensors, data analytics, and system integration make IoT-enabled CSCs difficult. Implementing, deploying, and administering these systems may be challenging without trained people. Integration of IoT into CSCs demands expertise. Without qualified workers, organizations may struggle to analyze this data.
[44] [45] [46] [47]
04
Data security and privacy
@@
Data generated and sent by IoT devices is often private. Extremely regulated sectors, such as healthcare and pharmaceuticals, make it difficult to safeguard this data and guarantee privacy compliance.
[48] [49] [36]
05
Technological infrastructure
@@
Integrating the IoT into an already established supply chain might be challenging. On the other hand, it’s likely that some of the more recent IoT technologies won’t be compatible with more established networks and software. IoT devices can’t be deployed and used effectively without the right infrastructure in place.
[31] [50] [36] [51]
06
Scalability challenges
@@
Scalability is a challenge for numerous sectors. It may be difficult to implement IoT throughout a complex supply chain. Cold supply networks need strict data monitoring and traceability. Implementing and running these systems need accurate data. Coordinating many linked devices and data becomes more difficult as IoT networks grow.
[52] [50] [53] [46]
07
Cultural and ethical consideration
@@
Cultural norms and expectations must be addressed when adopting IoT devices that monitor behaviors and data. IoT devices that monitor and enhance work processes may affect employees differently by culture. Technological developments may not benefit all workers. As IoT automation reduces human control, workers may fear they will be outdated.
[54] [40] [47]
08
Acceptance and adoption
@@
IoT implementation may involve organizational training and cultural changes; therefore, resistance to change and acceptance may be a significant obstacle. Several parties in the CSC may not understand IoT’s potential benefits, generating resistance. Worker resistance to the current system may hinder adoption.
[55] [56]
09
Data integration and management issue
@@
Constructing a cohesive and efficient system might be difficult due to possible communication difficulties across different IoT devices and platforms. The IoT is the source of a lot of data. Without proper data management and analytics tools, it is very challenging to deal with this massive amount of data, let alone analyze it and derive any meaningful conclusions from it.
[31] [57] [14] [53]
10
Reliability issue
@@
Inefficient asset monitoring and network concerns may cause asset loss, theft, customer distrust, delays, data loss, and inefficiency. IoT device failures may interrupt operations and raise maintenance expenses. Improper data input or modification without validation might affect reliability, resulting in non-compliance with industry standards and legal issues.
[28] [18] [42] [49]
11
ROI uncertainty
@@
A thorough cost-benefit analysis is required to determine the potential return on investment. The overall cost of an IoT deployment might be more than expected if the time it takes to see a return on investment is greater than expected.
[28] [29]
12
Extreme environment issue
@@
Extreme weather is a common occurrence for cold supply networks. Radio wave frequency is impacted by nearby elements, including metals and water, resulting in variable performance. It might be difficult to maintain the functionality and reliability of IoT devices when they are in cold storage or in transit.
[58] [28] [32] [59]
13
Cold chain network
@@
Remote or mobile supply chain equipment powering might be difficult. IoT devices may have trouble delivering data in rural or poor areas with poor network coverage. Many temperature-sensitive items can not resist slight temperature fluctuations during transport, storage, and handling, which could affect the cold chain network.
[60] [61] [62] [56] [59]
@@ 2.4 Methodology related literature
ISM is a computer-based learning method that assists individuals or groups in creating a visual representation of the intricate connections among several aspects of a complex situation. The fundamental idea is to use the practical skills and knowledge of specialists to dissect a complex system into many sub-systems and construct a multi-level structural model[9]. Raut et al. [63] employed the ISM approach to establish the interconnections among the drivers, enabling a comprehensive knowledge of the relative connections among the critical success factors but also defining their dependency when implementing sustainability.
The DEMATEL technique has been used proficiently for the purpose of identifying critical barriers in various vital processes, which was created to support methodical evaluations and organized decision-making processes. The effect connections map shows the relational linkages, with numbers denoting influence degree and arrows denoting influence direction. It is a comprehensive tool for assessing interrelations between system variables and emphasizing core-driving components [10]. Manoharan et al. [64] presented the DEMATEL model, which aids in comprehending the importance of drivers and barriers, creating matrices or digraphs to represent complex causal interactions, and discovering the relationships between these factors. Similar barrier analysis techniques employed in recent attempts are outlined in Table 2.
@@
1
[65]
ISM MICMAC
Figured out what matters most for reducing food waste and loss in supply chains for fresh produce
Regulatory bodies, food policy, and market infrastructure are the most important elements that have a strong propensity to affect the lowering of FLW in FPSCs.
2
[66]
DEMATEL-MMDE-ISM
Identified and evaluated potential barriers to IoT deployment in the manufacturing industry
Highlighted the key challenges where academics and practitioners may concentrate their strategic efforts to overcome implicit concerns while deploying IoT Techniques in manufacturing.
3
[67]
Bayesian Best-Worst Method
Discovered barriers to the implementation of the circular economy in the manufacturing sector.
Recommended a more realistic PFS performance, allowing several uses.
4
[68]
Bayesian Best-Worst Method
Identified multiple barriers to implementing the IoT in the food industry
Major barrier to the IoT in the food industry is the absence of an adequate internet infrastructure.
5
[69]
ISM-DEMATEL
Examined the identified obstacles that impede the implementation of electronic waste management practices in India.
Key challenges pertaining to e-waste recycling are a lack of public knowledge and rules about the matter.
6
[70]
DEMATEL–ISM
Identified what circumstances affected the flight crew’s TSA protocol
Different levels of hierarchy, mutual impact, and the cause-and-effect link are shown for the flight crew’s TSA creation and management.
7
[71]
Fuzzy-ISM-DEMATEL
Identified barriers to implementing sustainable purchasing practices over the internet
Improved DEMATEL, ISM, and rough set theory into a unified framework.
8
[72]
ISM-DEMATEL
Analysed the challenges associated with waste recycling in India
For India’s trash recycling infrastructure to grow, significant barriers such as a lack of funds, input material, and subsidies have to be overcome.
The literature investigation talks about every element of the CSC, covering transportation, cold storage, procurement, and occupational security across the cold chain. IoT supports SCM, enhancing visibility and stability, but problems include regulatory limitations and bad infrastructure. Most of the study focuses on distinct problems and seeks to solve these challenges with the provided solutions. In the combined ISM-MICMAC-DEMATEL approach, ISM-MICMAC is used for constructing hierarchical relationships among barriers and analyzing the structure of systems, while DEMATEL delivers a numerical evaluation of the strength and direction of interactions between factors. Numerous barrier analysis strategies are employed in modern methods that are also used to analyze obstacles. The choice between these tactics varies depending on individual circumstances, available data, and decision-makers’ preferences. From thirty-one research papers, thirteen barriers were found, and from the different barrier analysis techniques utilized in recent approaches, ISM-DEMATEL was chosen to analyze these barriers.
@@ 3 Methodology
The research initiation involves several key stages, including the design of surveys, meticulous data collection, selection procedures, and rigorous reliability testing, as illustrated in Fig 2. Subsequently, the second phase unfolds with a process flow diagram outlining the development of an ISM model. The third phase is characterized by a process flow diagram that delineates the systematic steps employed in constructing the DEMATEL model.
@@ 3.1.1 Barriers identification
This involves identifying and recognizing possible obstacles or challenges inside the system. These obstacles might be technical, organizational, or connected to other system operation elements. With the use of literature research and expert comments, potential barriers to the system are identified, as described in Table 1.
@@ 3.1.2 Survey development and data selection procedure
Thirty experts participated in the ISM survey. The diagonal values in the ISM method are all zero. Only use one side of the diagonal value to create the initial reachability matrix. With the exception of the diagonal values, all thirteen rows and thirteen columns of the matrix were surveyed. This verification process guarantees that the data provided by the respondents is valid. After determining that the impact of barrier 1 on barrier 2 is V, the expert must have concluded that the impact of barrier 2 on barrier 1 is A. If O and X are on a diagonal, then their opposing sides must be the same. These checks will improve the survey’s quality, even if the initial reachability matrix ultimately decides the whole system.
The DEMATEL impact factor survey, which is likewise associated with the ISM survey, is also measured on a scale from 0 to 4. The next issue concerns the impact factor, which must be larger than or equal to 1 if the effect of barrier 1 on barrier 2 is V. A value of 0 is required for the influence factor if it is O, and a positive number is also required for X. Therefore, this method of cross-checking will enhance the survey’s quality.
Twelve reasonable answers were selected for each question, and twelve $13\times 13$ matrices were constructed from them. The average of these matrices was then calculated to determine the average matrix. For ISM, the structural self-interaction matrix is also prepared using the value that appears most often in these logical replies.
Table 3 provides a succinct overview of the varied characteristics present in the expert panel. The quantification of each category offers vital information about the group’s makeup. A table of this kind assists in visually representing and examining the distribution of expertise. The organized survey data from these thirty experts helps the comprehensibility of the information, which is essential for generating the outcome. The gender distribution comprises 26 males and 4 females. Regarding age, 13 panelists are below 30, 8 fall in the 30–40 range, 5 are aged 41–50, and 4 are 51 years and above. Regarding experience in their respective fields, 10 experts have less than five years, 7 possess 5–10 years, 10 have 10–15 years, and 6 hold 15 years and above. The domain of expertise includes 8 in Academics (Industrial Engineering), 5 in System Integration, 4 in Data Analysts, 6 in Data Security, and 7 in Cold Chain Executives.
@@ 3.1.3 Establish relationships between barriers
Once the possible obstacles are discovered, it’s essential to understand how these barriers are interrelated or impact each other. This step assists in understanding the system’s complexity and how different challenges may be connected. In order to determine which pairs of barriers need further investigation, the contextual relationship between them is established.
@@ 3.1.4 Cronbach’s alpha testing
The reliability coefficient Cronbach’s alpha can be used for evaluating survey or test item internal consistency. Many psychological and educational surveys and assessments use Cronbach’s alpha in their formulation and testing. It aids researchers and practitioners in ensuring their measuring equipment accurately measures the idea in question. Cronbach’s alpha reliability is a measure of how well a sum or average of k measurements holds up, where k measures might be anything from k raters to alternate forms to items on a questionnaire or exam. We have a 169-item questionnaire (k = 156) to measure implementation barrier influence, with each item scored on a 0–4 Likert scale. A Cronbach’s alpha value should exceed 0.7 since a value equal to or more than 0.7 falls within an acceptable range [73].
The variance of the total score (270.52) and the sum of the item variances (67.88) provide this dependability measure. The survey dataset has an acceptable degree of internal consistency, as shown by the Cronbach’s alpha value of 0.7534. This measuring tool reliably captures the targeted construct since the coefficient is greater than the generally recognized requirements of 0.7.
@@ 3.2 Interpretative structural model (ISM)
The initial development of ISM can be attributed to J. Warfield in 1973 to examine intricate socioeconomic systems [72]. It is a method that may help individuals or teams better grasp a topic by organizing their understanding of it into a network diagram. The following is a detailed description of how ISM works [11][12].
@@ 3.2.1 Step 1: Developing structural self-interaction matrix (SSIM)
The pair-wise connection between barriers in the system can be observed by employing SSIM for barriers. The following four symbols each indicate a unique set of relationships:

### v4

@@ 2.3 IoT implementation barriers
IoT deployment is challenging because of device and data security, interoperability issues, scalability, trust, privacy concerns, financial issues, lack of qualified personnel, and compliance with data privacy laws. First, it is necessary to identify the barriers that hinder IoT implementation in CSC in order to overcome these challenges. Table 1 lists the barriers to implementing the IoT in CSC.
@@
01 Regulatory compliance The CSC has strict temperature and data management guidelines. It may be challenging to implement IoT technology and meet these requirements. In the pharmaceutical and food industries, strict regulations are tough to follow. Failing to satisfy these regulations may result in product recalls or legal concerns. [52] [53] [54] [55] [48]
@@
02 High implementation cost In order to adopt the IoT, it may be necessary to make a sizable initial investment in things like sensors, devices, and infrastructure. Transporting and storing cold requires plenty of energy. Costs may rise with power prices. [56] [39] [40] [51]
@@
03 Lack of skilled workforce Multiple sensors, data analytics, and system integration make IoT-enabled CSCs difficult. Implementing, deploying, and administering these systems may be challenging without trained people. Integration of IoT into CSCs demands expertise. Without qualified workers, organizations may struggle to analyze this data. [57] [58] [59] [60]
@@
04 Data security and privacy Data generated and sent by IoT devices is often private. Extremely regulated sectors, such as healthcare and pharmaceuticals, make it difficult to safeguard this data and guarantee privacy compliance. [61] [62] [47]
@@
05 Technological infrastructure Integrating the IoT into an already established supply chain might be challenging. On the other hand, it’s likely that some of the more recent IoT technologies won’t be compatible with more established networks and software. IoT devices can’t be deployed and used effectively without the right infrastructure in place. [42] [63] [47] [64]
@@
06 Scalability challenges Scalability is a challenge for numerous sectors. It may be difficult to implement IoT throughout a complex supply chain. Cold supply networks need strict data monitoring and traceability. Implementing and running these systems need accurate data. Coordinating many linked devices and data becomes more difficult as IoT networks grow. [65] [63] [66] [59]
@@
07 Cultural and ethical consideration Cultural norms and expectations must be addressed when adopting IoT devices that monitor behaviors and data. IoT devices that monitor and enhance work processes may affect employees differently by culture. Technological developments may not benefit all workers. As IoT automation reduces human control, workers may fear they will be outdated. [67] [53] [60]
@@
08 Acceptance and adoption IoT implementation may involve organizational training and cultural changes; therefore, resistance to change and acceptance may be a significant obstacle. Several parties in the CSC may not understand IoT’s potential benefits, generating resistance. Worker resistance to the current system may hinder adoption. [68] [69]
@@
09 Data integration and management issue Constructing a cohesive and efficient system might be difficult due to possible communication difficulties across different IoT devices and platforms. The IoT is the source of a lot of data. Without proper data management and analytics tools, it is very challenging to deal with this massive amount of data, let alone analyze it and derive any meaningful conclusions from it. [42] [70] [16] [66]
@@
10 Reliability issue Inefficient asset monitoring and network concerns may cause asset loss, theft, customer distrust, delays, data loss, and inefficiency. IoT device failures may interrupt operations and raise maintenance expenses. Improper data input or modification without validation might affect reliability, resulting in non-compliance with industry standards and legal issues. [39] [22] [55] [62]
@@
11 ROI uncertainty A thorough cost-benefit analysis is required to determine the potential return on investment. The overall cost of an IoT deployment might be more than expected if the time it takes to see a return on investment is greater than expected. [39] [40] [49]
@@
12 Extreme environment issue Extreme weather is a common occurrence for cold supply networks. Radio wave frequency is impacted by nearby elements, including metals and water, resulting in variable performance. It might be difficult to maintain the functionality and reliability of IoT devices when they are in cold storage or in transit. [71] [39] [43] [38]
@@
13 Cold chain network Powering remote or mobile supply chain equipment might be difficult. IoT devices may have trouble delivering data in rural or poor areas with poor network coverage. Many temperature-sensitive items can not resist slight temperature fluctuations during transport, storage, and handling, which could affect the cold chain network. [72] [73] [74] [69] [38]
@@ 2.4 Methodology related literature
ISM is a computer-based learning method that assists individuals or groups in creating a visual representation of intricate connections between several aspects of a complex situation. The fundamental idea is to use the practical skills and knowledge of specialists to dissect a complex system into many sub-systems and construct a multi-level structural model[11]. Raut et al. [75] employed the ISM approach to establish the interconnections among drivers, enabling a comprehensive knowledge of the relative connections among the critical success factors and defining their dependency when implementing sustainability.
The DEMATEL technique has been used extensively to identify critical barriers in various vital processes, which were created to support methodical evaluations and organized decision-making processes. The effect connections map shows the relational linkages, with numbers denoting the degree of influence and arrows denoting the direction of influence. It is a comprehensive tool for assessing the interrelations between system variables and emphasizing the core-driving components [12]. Manoharan et al. [76] presented the DEMATEL model, which aids in comprehending the importance of drivers and barriers, creating matrices or digraphs to represent complex causal interactions, and identifying the relationships between these factors. The barrier analysis techniques employed in recent studies are presented in Table 2.
@@
1 [77] ISM MICMAC Figured out what matters most for reducing food waste and loss in supply chains for fresh produce Regulatory bodies, food policy, and market infrastructure are the most important elements that have a strong propensity to affect the lowering of FLW in FPSCs.
2 [78] DEMATEL-MMDE-ISM Identified and evaluated potential barriers to IoT deployment in the manufacturing industry Highlighted the key challenges where academics and practitioners may concentrate their strategic efforts to overcome implicit concerns while deploying IoT Techniques in manufacturing.
3 [79] Bayesian Best-Worst Method Discovered barriers to implementing the circular economy in the manufacturing sector. Recommended a more realistic PFS performance, allowing several uses.
4 [80] Bayesian Best-Worst Method Identified multiple barriers to implementing the IoT in the food industry Major barrier to the IoT in the food industry is the absence of an adequate internet infrastructure.
5 [81] ISM-DEMATEL Examined the identified obstacles that impede the implementation of electronic waste management practices in India. Key challenges pertaining to e-waste recycling are a lack of public knowledge and rules about the matter.
6 [82] DEMATEL–ISM Identified what circumstances affected the flight crew’s TSA protocol Different levels of hierarchy, mutual impact, and the cause-and-effect link are shown for the flight crew’s TSA creation and management.
7 [83] Fuzzy-ISM-DEMATEL Identified barriers to implementing sustainable purchasing practices over the internet Improved DEMATEL, ISM, and rough set theory into a unified framework.
8 [84] ISM-DEMATEL Analysed the challenges associated with waste recycling in India For India’s trash recycling infrastructure to grow, significant barriers such as a lack of funds, input material, and subsidies have to be overcome.
The literature investigation discusses every element of the CSC, covering transportation, cold storage, procurement, and occupational security across the cold chain. IoT supports SCM by enhancing visibility and stability, but problems include regulatory limitations and poor infrastructure. Most studies focus on distinct problems and seek to solve these challenges using the solutions provided. In the combined ISM-MICMAC-DEMATEL approach, ISM-MICMAC is used to construct hierarchical relationships among barriers and analyze the structure of systems, while DEMATEL delivers a numerical evaluation of the strength and direction of interactions between factors. Numerous barrier analysis strategies have been employed in modern methods, which are also used to analyze obstacles. The choice between these tactics varies depending on individual circumstances, available data, and decision-makers’ preferences. Thirteen barriers, and from the different barrier analysis techniques utilized in recent approaches, ISM-DEMATEL was chosen to analyze these barriers.
@@ 3 Methodology
Research initiation involves several key stages, including the design of surveys, meticulous data collection, selection procedures, and rigorous reliability testing, as illustrated in Fig 2. Subsequently, the second phase unfolds with a process flow diagram outlining the development of the ISM model. The third phase is characterized by a process flow diagram that delineates the systematic steps employed in constructing the DEMATEL model.
@@ 3.1.1 Barriers identification
This involves identifying and recognizing the possible obstacles or challenges inside the system. These obstacles may be technical, organizational, or connected to other system operation elements. Using literature research and expert comments, potential barriers to the system were identified, as described in Table 1.
@@ 3.1.2 Survey development and data selection procedure
Thirty experts participated in the survey. The diagonal values in the ISM method are zero. Only one side of the diagonal value is used to create the initial reachability matrix. All 13 rows and 13 columns of the matrix were surveyed, with the exception of diagonal values. This verification process guaranteed that the data provided by the respondents were valid. After determining that the impact of barrier 1 on barrier 2 is V, an expert must conclude that the impact of barrier 2 on barrier 1 is A. If O and X are diagonal, their opposing sides must be the same. These checks will improve the quality of the survey, even if the initial reachability matrix ultimately determines the entire system.
The DEMATEL impact factor survey, which is likewise associated with the ISM survey, is also measured on a scale from 0 to 4. The next issue concerns the impact factor, which must be greater than or equal to 1 if the effect of barrier 1 on barrier 2 is V. A value of 0 is required for the influence factor if it is O. A positive number is also required for X. Therefore, this cross-checking method will enhance the quality of the survey.
Twelve reasonable answers were selected for each question, and twelve $13\times 13$ matrices were constructed. The average of these matrices was calculated to determine the average matrix. For the ISM, the structural self-interaction matrix is also prepared using the value that appears most often in these logical replies.
Table 3 provides a brief overview of the various characteristics of the expert panel. Quantification of each category offers vital information regarding the makeup of the group. A table of this type assists in visually representing and examining the distribution of expertise. The organized survey data from these 30 experts helps the comprehensibility of the information, which is essential for generating the outcome. The sex distribution comprised 26 males and 4 females. Thirteen panelists are below 30, 8 fall in the 30–40 range, 5 are aged 41–50, and 4 are 51 years and above. Regarding experience in their respective fields, 10 experts had less than five years, 7 possessed 5–10 years, 10 had 10–15 years, and 6 held 15 years and above. The domain of expertise includes eight in academics (Industrial Engineering), five in System Integration, four in Data Analysts, six in Data Security, and seven in cold chain executives.
The survey was conducted using Google Forms, and the recruitment period was from 27 Sep 2023 to 11 Oct 2023. Participants were selected based on their expertise in supply chain, system integration, data analysis, data security, and cold chain. The survey consisted of 312 questions, focusing on the influential relationships among the barriers. Confidentiality of participant responses was strictly maintained. All the collected data were anonymized and stored. Only the research team had access to the raw data, and the findings were reported in aggregate form to ensure the anonymity of the participants. This study adhered to the ethical principles outlined in the Declaration of Helsinki. Participants were provided with informed consent forms detailing the purpose of the study, their rights as participants, and the procedures for data handling. Participation in the survey was voluntary, and participants had the right to withdraw at any time without penalties. This study was approved by the Ethics Review Board of the Research & Extension Center, Khulna University of Engineering & Technology, Bangladesh, to conduct this survey research.
@@ 3.1.3 Establish relationships between barriers
Once possible obstacles are discovered, it is essential to understand how these barriers are interrelated or impact each other. This step helps in understanding the complexity of the system and how different challenges may be connected. To determine which pairs of barriers require further investigation, a contextual relationship between them is established.
@@ 3.1.4 Cronbach’s alpha testing
The reliability coefficient Cronbach’s alpha can be used to evaluate the internal consistency of the survey or test items. Many psychological and educational surveys and assessments have used Cronbach’s alpha for their formulation and testing. This aids researchers and practitioners in ensuring that their measuring equipment accurately measures the idea in question. Cronbach’s alpha reliability is a measure of how well a sum or average of k measurements holds, where k measures might be anything from k raters to alternate forms to items on a questionnaire or exam. We used a 156-item questionnaire (k = 156) to measure implementation barrier influence, with each item scored on a 0–4 Likert scale. A Cronbach’s alpha value should exceed 0.7 since a value equal to or more than 0.7 falls within an acceptable range [85].
The variance of the total score (270.52) and the sum of the item variances (67.88) provide this dependability measure. The survey dataset had an acceptable degree of internal consistency, as indicated by a Cronbach’s alpha value of 0.7534. This measuring tool reliably captures the targeted construct because the coefficient is greater than the generally recognized requirement of 0.7.
@@ 3.2 Interpretative structural model (ISM)
The initial development of ISM can be attributed to J. Warfield in 1973 who examined intricate socioeconomic systems [84]. This method may help individuals or teams better grasp a topic by organizing their understanding of it into a network diagram. The following is a detailed description of how ISM works [13][14].
@@ 3.2.1 Step 1: Developing structural self-interaction matrix (SSIM)
The pair-wise connection between barriers in the system can be observed by employing SSIM. The following four symbols indicate a unique set of relationships: